OpenStudy (anonymous):

Which of the following is greatest? seven twenty-fifths zero point three three nineteen one hundredths eleven-fiftieths

OpenStudy (lgbasallote):

\[\frac{7}{25}\] \[0.33 \implies \frac{1}{3}\] \[\frac{19}{100}\] \[\frac{11}{50}\]

OpenStudy (lgbasallote):

which one do you think?

OpenStudy (anonymous):

lol geez just say the answer

OpenStudy (anonymous):

7/25=0.28 0.33 19/100=0.19 11/50=0.22 so 0.33 is grtst

OpenStudy (anonymous):

arch1 can you you help me with a math problemm plz

OpenStudy (anonymous):

can some one help me plz if u like message me

OpenStudy (anonymous):

@wilnie here! :D

OpenStudy (anonymous):

@wilnie: ill try